Two cars leave town at the same time going in opposite directions. One travels 55 mph and the other travels 48 mph. In how many hours will they be 206 miles apart?
Answer by checkley77(12844)

55T+48T=206
103T=206
T=206/103
T=2 HOURS THEY WILL BE 206 MILES APART.
pROOF:
55*2+48*2=206
110+96=206
206=206